# Enno Adler    Stefan Böttcher    Rita Hartel

### v2

@@ Introduction
We investigate the time needed to answer a neighborhood query or a triple query. We compare results of the compressors listed in Table 1 (a). We test all implementations on a Debian 5.10.149-2 machine with 128GB RAM and 2 Cores @ 2.30GHz. For the tests, we use the RDF graphs homepages-en, geo-coorindates-en, jamendo, wikidata, archiveshub, scholarydata-dump as in Röder et al. [6], the version graphs ttt-win and chess-legal from SUBDUE [^1] https://ailab.wsu.edu/subdue/download.htm and the webgraphs CA-AstroPh, NotreDame and WikiTalk from the Standford Large Network Dataset Collection [14]. As not all compressors are designed to use webgraphs and version-graphs as input, we converted their structure to an RDF notation and take these converted files as input for all approaches for fairness.
ITR is implemented [^2] The implementation of ITR is available at https://github.com/adlerenno/IncidenceTypeRePair in C. For the $k^{2}$-tree approach, we use the implementations of Röder et al. [6] as in their paper in Java and C++, $k^{2}$-java and $k^{2}$++, respectively. We include two implementations of HDT, HDT-java and HDT++.
In Figure 3, we see that ITR compresses stronger than the existing solutions on some datasets like geo-coordinates, and archiveshub, but other approaches compress stronger on the network and version graphs due to a non-optimal dictionary compression.
@@
Due to the different supported query types, we compare the runtime by using queries that are equal if possible. For example, we use SELECT ?pr ?ob WHERE { v ?pr ?ob } as SPARQL query and v ?P ?O as triple query for an outgoing neighborhood query of the node $v$. As the gRePair implementation does not support queries with labels, gRePair is only included in the comparisons of S ? ? and ? ? O. We perform 500 queries of each query type on each file and show the average time needed to answer the queries.
Figure 3: The compression ratio is the file size of the compressed graph divided by the file size of the uncompressed input file. We stopped RDFRePair on wikidata after 6 days.
Figure 4: Average runtime in milliseconds of 500 queries each on the geo-coordinates-en graph. Implementations that do not appear in the legend do not support any type of query. gRePair can only answer neighborhood queries (S ? ? and ? ? O).
In Figure 4, we see that gRePair, HDT-java, and $k^{2}$-java as Java or Scala implementations are outperformed by the C and C++ implementations. ITR is about 2 to 6 times faster than HDT++ except for ? P ? queries. Furthermore, ITR is more than 100 times faster than gRePair.
Summary and Conclusion
We presented the ITR graph compressor based on a new type of digrams and index-functions. Futhermore, ITR+ can use hyperedges of rank 1 for frequent node labels to reduce the needed dictionary size to store node labels. Our evaluation of the runtime of triple queries shows that ITR performs all queries except ? P ? significantly faster than the existing variants of RePair on graphs, namely gRePair and RDFRePair, and faster than $k^{2}$-tree and HDT provided on different implementations.

### v3

@@ 4 Experimental Results
We investigate the time needed to answer a neighborhood query or a triple query. We compare the results of the compressors listed in Table 1 (a). We test all implementations on a Debian 5.10.209-2 machine with 128GB RAM and 32 Cores Intel(R) Xeon(R) Platinum 8462Y+ @ 2.80GHz. For the tests, we use the RDF graphs homepages-en, geo-coorindates-en, jamendo, wikidata, archiveshub, scholarydata-dump as in Röder et al. [6], the version graphs ttt-win and chess-legal from SUBDUE [^2] https://ailab.wsu.edu/subdue/download.htm, and the webgraphs CA-AstroPh, NotreDame, and WikiTalk from the Standford Large Network Dataset Collection [13].
ITR is implemented [^3] The implementation of ITR is available at https://github.com/adlerenno/IncidenceTypeRePair in C. For the $k^{2}$-tree approach, we use the implementations of Röder et al. [6] as in their paper in Java and C++, $k^{2}$-java and $k^{2}$++, respectively. We include two implementations of HDT, namely HDT-java [7] and HDT++ [10]. We omit RDF engines because RDF engines mostly cannot handle multiple nodes with the same label and thus, RDF is only a subclass of graphs we look at.
In Figure 3, we see that ITR compresses better than the existing solutions on some datasets like geo-coordinates, and archiveshub, but other approaches compress better on the network and version graphs due to a non-optimal dictionary compression.
Figure 3: The compression ratio is the file size of the compressed graph divided by the file size of the uncompressed input file. We stopped RDFRePair on wikidata after 6 days. For ITR+, only chess-legal and ttt-win use the same node label for multiple nodes.
Figure 4: Average runtime in milliseconds of 500 queries each on the jamendo graph. Implementations that do not appear in the legend do not support any type of query. gRePair can only answer neighborhood queries (S ? ? and ? ? O).
@@
Due to the different supported query types, we compare the runtime by using queries that are equal. As the gRePair implementation does not support queries with edge labels, gRePair is only included in the comparisons of S ? ? and ? ? O. We use SELECT ?pr ?ob WHERE { v ?pr ?ob } as SPARQL query and v ?P ?O as triple query for an outgoing neighborhood query of the node $v$. We perform 500 queries of each query type on each file and show the average time needed to answer the queries.
In Figure 4, we see that gRePair, HDT-java, and $k^{2}$-java as Java or Scala implementations are outperformed by the C and C++ implementations. ITR is about 2 to 6 times faster than HDT++ except for ? P ? queries. Furthermore, ITR is more than 100 times faster than gRePair. In Figure 4, we presented the results for the jamendo graph, but the time ratios of ITR to other approaches are similar for all tested graphs.
5 Summary and Conclusion
We presented the ITR graph compressor based on a new type of digrams and on index-functions. Furthermore, ITR+ can use hyperedges of rank 1 for frequent node labels to reduce the needed dictionary size to store node labels. Our evaluation of the runtime of triple queries shows that ITR performs all queries except ? P ? significantly faster than the existing variants of RePair on graphs, namely gRePair and RDFRePair, and faster than $k^{2}$-tree and HDT provided on different implementations.
@@ Acknowledgements
Appendix A: On the Accuracy of the Count Estimate
In Appendix A, we investigate the accuracy of the presented counting of digrams. Let $G$ be a fixed graph and $d=((a,k),(b,l))$ be a fixed digram. Let $O$ be a set of non-overlapping occurrences of $d$ and $O$ be maximal in size. We show that $count(d)\geq|O|$, and $count(d)=|O|$ if either $a\neq b$ or $k=l$.
We define $O_{v}=\{(e_{1},e_{2})\in O|e_{1}[k]=e_{2}[l]=v\}$ as the subset of $O$ that has the node $v$ as shared node defined by the incidence-types of digram $d$. For two nodes different nodes $v_{1}$ and $v_{2}$, $O_{v_{1}}\cap{}O_{v_{2}}=\emptyset$, because we get $v_{1}=e_{1}[k]=v_{2}$ for all $e_{1}\in O_{v_{1}}\cap{}O_{v_{2}}$ as an obvious contradiction otherwise. Hence each occurrence $(e_{1},e_{2})\in O$ is in $O_{e_{1}[k]}$ ($=O_{e_{2}[l]}$), the sets $O_{v}$ for all nodes $v\in V$ form a disjoint union of $O$, so $O=\mathop{\vphantom{\bigcup}\mathchoice{\leavevmode\vtop{\halign{\hfil$\m@th\displaystyle#$\hfil\cr\bigcup\cr\cdot\crcr}}}{\leavevmode\vtop{\halign{\hfil$\m@th\textstyle#$\hfil\cr\bigcup\cr\cdot\crcr}}}{\leavevmode\vtop{\halign{\hfil$\m@th\scriptstyle#$\hfil\cr\bigcup\cr\cdot\crcr}}}{\leavevmode\vtop{\halign{\hfil$\m@th\scriptscriptstyle#$\hfil\cr\bigcup\cr\cdot\crcr}}}}\displaylimits_{v\in V}O_{v}$.
Regarding a single fixed node $v$, $E_{1}=\{e\in E|label(e)=a\text{ and }e[k]=v\}$ and $E_{2}=\{e\in E|label(e)=b\text{ and }e[l]=v\}$ contain possible edges for the first and second incidence-type of the digram $d$, respectively. If the two incidence-types are equal, $E_{1}=E_{2}$ and $\left\lfloor\frac{|E_{1}|}{2}\right\rfloor$ is the maximum number of pairs that can be drawn from $E_{1}$ without repetition. If the incidence-types are different, $E_{1}\neq E_{2}$ and consequently $min(|E_{1}|,|E_{2}|)$ is the maximum number of non-overlapping pairs of edges. We define $count_{v}$ accordingly.
[…]
$=\sum_{v\in V}|O_{v}|$
$\leq\sum_{v\in V}count_{v}(d)$
$=count(d)$
This shows that our estimate of the number of digram occurrences is an upper bound. Assuming $a\neq b$ or $k=l$, each edge $e\in E$ with label $a$ (or $b$) belongs to exactly one node $e[k]$ (or $e[l]$). Only at $e[k]$ (or $e[l]$) it is possible to form a digram using the edge $e$. $|O_{v}|=count_{v}(d)$ follows.
Only the case $a=b$ and $k\neq l$ violates in two ways our counting: Overlapping occurrences with different shared nodes or loops with the incidence-types of the digram are not counted correctly. An completely accurate count would take $\mathcal{O}(|E|^{2})$ [5], which is infeasible.